# Qiuyu Wang2,    Hao Ouyang2,    Hao Chen1,    Chunhua Shen1,3,2

### v1

@@ 4.1 Temporal Self-Consistency Voting
In this section, we propose a temporal self-consistency decoding strategy for dLLMs, which leverages intermediate predictions during sampling to produce more reliable final outputs.
As discussed in Sec. 3.3, although the final timestep in a diffusion trajectory typically yields the most accurate results overall, the correct answer to a given question may sometimes appear earlier in the trajectory. Relying solely on the final prediction thus risks discarding better intermediate outputs. To address this, we propose a method that aggregates predictions across timesteps through a weighted voting mechanism.
Formally, given a diffusion sampling trajectory $\left\{x_{0}^{t}\right\}_{t=1}^{T}$, our method selects the final answer $a^{*}$ according to a weighted vote over all timesteps:
@@
Here, $\mathbbm{1}(\cdot)$ is the indicator function that returns 1 if the decoded meaning of $x_{0}^{t}$ matches the candidate answer $a$, and $f(t)$ is a weighting function over timesteps. Since the accuracy of predictions generally tends to improve as the sampling step increases (or diffusion step decreases), we design $f(t)$ to be a monotonically decreasing function of diffusion step $t$. In experiments, we explore three types of weighting schemes: constant, linear decay, and exponential decay, as detailed in Sec. 5.2.
Discussion. Our method is conceptually related to self-consistency decoding (Wang et al., 2022), which improves reasoning in autoregressive LLMs by sampling diverse reasoning paths and selecting the most consistent answer via majority voting. However, self-consistency requires multiple full-length forward passes, leading to significant computational overhead. In contrast, our approach requires only a single sampling trajectory from the dLLMs. By exploiting the temporal nature of diffusion inference indicated by Eq. 1, we obtain a series of intermediate predictions essentially for free, without additional model evaluations. This makes our method both efficient and effective for boosting accuracy through temporal aggregation.
@@ 4.2 Temporal Consistency Reinforcement
Negative TSE as the Reward. Following GRPO, for each question $q$ sampled from the dataset $\mathcal{D}$, we draw a group of $G$ responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the current policy $\pi_{\theta}$. Each response $o_{i}$ receives a scalar reward $r_{i}=-\operatorname{TSE}(o_{i})$, where $\operatorname{TSE}(o_{i})$ is computed using Eq. 3 from Sec. 3.3. This reward encourages the model to produce responses whose intermediate predictions remain semantically consistent throughout the decoding process. Based on this, we define the unnormalized advantage (Liu et al., 2025b) for all tokens $k=1,\ldots,|o_{i}|$ as $A_{i}^{k}(\pi)=r_{i}(\pi)-\operatorname{mean}\left(\left\{r_{j}(\pi)\right\}_{j=1}^{G}\right).$ The training objective follows the standard GRPO formulation with our TSE-based reward:
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{\stackrel{{\scriptstyle q\sim\mathcal{D}}}{{o_{1},\ldots,o_{G}\sim\pi_{\theta}(\cdot|q)}}}\left[\left(\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\min\left(\rho_{i}^{k}A_{i}^{k},\text{clip}\left(\rho_{i}^{k},1-\varepsilon,1+\varepsilon\right)A_{i}^{k}\right)\right)-\beta D_{\text{KL}}\left[\pi_{\theta}(\cdot|q)\|\pi_{\text{ref}}(\cdot|q)\right]\right],$
@@
Combining TSE with Accuracy Reward. As an alternative, we can combine TSE with the accuracy reward to achieve further performance gain when ground-truth answers are available during training. We draw on insights from statistical decision theory, specifically the framework of proper scoring rules (Gneiting & Raftery, 2007; Damani et al., 2025). These scoring rules are designed to encourage honest confidence estimates: they are minimized when the predicted confidence $c$ accurately reflects the true likelihood that the model’s output $o$ matches the correct answer $o^{*}$.
We derive the confidence $c$ from the TSE by normalizing it with respect to the maximum entropy achievable under uniform sampling. Specifically, we define confidence as $c(o_{i})=\frac{\mathcal{H}_{\max}-\operatorname{TSE}(o_{i})}{\mathcal{H}_{\max}}$, where $\mathcal{H}_{\max}=\log T$. Here, $T$ denotes the total number of sampling steps. The normalization ensures $c\in[0,1]$, where higher values indicate greater confidence for correct answers.
The accuracy reward follows a binary scheme: it assigns a reward of 1 when the model’s prediction is correct (i.e., $o_{i}=o^{*}$), and 0 otherwise. We propose a composite reward function grounded in the theory of proper scoring rules. Specifically, we adopt the spherical scoring rule, which empirically demonstrates superior performance across tasks. The final reward function is defined as:
@@
Here, the second term incorporates the model’s confidence, rewarding high certainty when the prediction is correct while penalizing overconfidence in incorrect predictions. This design ensures the reward is sensitive not just to accuracy, but also to the reliability of the prediction. A detailed comparison of various scoring rules and their empirical performance in the Sec. D.1.
@@ 5 Experiments
Table 1: Performance of temporal majority voting on mathematics benchmarks: We evaluate three temporal majority voting strategies, including fixed weighting, linear weighting, and exponential weighting, across four math datasets and two model variants: LLaDA-8B-Instruct and LLaDA-1.5. Exponential weighting achieves the best results across datasets and sequence lengths. Bold numbers indicate the highest performance within each group, while green values represent improvements over the baseline. For reference, we include the oracle $\operatorname{EverPass}@1\mid t$ as a performance upper bound.
@@ 5.1 Implementation Details
We utilize both LLaDA-8B-Instruct (Nie et al., 2025) and LLaDA-1.5 (Zhu et al., 2025), and evaluate performances on four widely recognized mathematical datasets: GSM8K (Cobbe et al., 2021), MATH500 (Hendrycks et al., 2021), SVAMP (Patel et al., 2021), and Countdown (Pan et al., 2025). In line with d1 (Zhao et al., 2025), we assess the model’s performance across various output lengths. For the temporal self-consistency voting, we employ the exponential function by default to weight the sampling steps. In terms of post-training, we follow d1 and perform supervised fine-tuning (SFT) before reinforcement fine-tuning (RFT) for LLaDA-8B-Instruct. Specifically, we conduct SFT on the s1K dataset (Muennighoff et al., 2025) for 20 epochs, using sequences with a length of 4,096. For LLaDA-1.5, we omit SFT because it often results in performance degradation, likely due to the model having already undergone sophisticated post-training. All training is executed using 8 H800 GPUs. Please refer to the Appendix B for more details.
@@ 5.2 Temporal Self-Consistency Voting
Voting Strategies. In experiments, temporal voting is performed by applying weighted voting across denoising steps using three schemes: fixed, linear, and exponential. Each uses a weighting function $f(t)$, where $t$ is the current diffusion step. The fixed scheme assigns equal weight to all steps with $f(t)=1$. The linear scheme emphasizes earlier diffusion steps using f(t)=1t/Tf(t)=1-t/T, while the exponential scheme takes $f(t)=\exp(\alpha(1-t/T))$, where $\alpha=5$ by default. Both linear and exponential schemes prioritize early diffusion time steps, i.e., latter sampling steps.
Ablations on Varying Voting Strategies. As shown in Table 1, both linear and exponential weighting improve the model’s inference performance. For example, exponential weighting consistently enhances the performance of LLaDA-8B-Instruct across different datasets, with an average improvement of 1.6%, 1.2%, 1.0% and 2.2% accuracy on the GSM8K, MATH500, SVAMP, and Countdown datasets, respectively. The model performs slightly worse under the fixed weighting strategy, which we attribute to the inaccuracy of the answers predicted in the early sampling steps; assigning the same weight to all steps leads to performance degradation. By default, we adopt the exponential weighting strategy.
We further conducted an ablation study on the hyperparameter $\alpha$ in the exponential weighting. Taking LLaDA-8B-Instruct as an example, we calculated the average improvement across four datasets and various response lengths. As illustrated in Fig. 5b, $\alpha$ values ranging from 1 to 11 consistently yielded improvements. The optimal performance was observed at $\alpha=5$, achieving an average improvement of 1.5% in accuracy. Based on these results, we set $\alpha=5$ as the default value.
@@ 5.3 Temporal Consistency Reinforcement
Table 2: Performance of reinforcement fine-tuning on mathematics benchmarks. Unlike the competing method d1 (Zhao et al., 2025), which requires access to ground-truth answers during training, our approach leverages the negative Temporal Semantic Entropy (TSE) reward without relying on ground-truth labels. Furthermore, combining TSE with accuracy-based rewards leads to even greater performance improvements across multiple benchmarks. The LLaDA baseline is obtained by applying supervised fine-tuning to LLaDA-8B-Instruct using the s1K dataset (Sahoo et al., 2024). Green numbers show the difference between our method and the baseline.
@@
+ RFT accuracy reward (d1) 70.6 78.5 81.4 29.6 34.2 39.2 85.0 88.0 88.3 32.8 25.8 39.5
negative TSE reward (ours) 70.3 79.5 81.4 29.2 35.6 39.0 86.0 88.0 88.7 34.0 48.8 49.5
+0.5 +0.1 +0.3 +0.2 +3.2 +3.6 +0.7 +1.7 +5.4 +12.5 +27.7 +28.8
combining both (ours) 70.2 79.5 81.7 30.8 34.8 41.2 88.7 89.3 89.0 53.1 43.0 55.1
+0.4 +0.1 +0.6 +1.8 +2.4 +5.8 +3.4 +3.0 +5.7 +31.6 +21.9 +34.4
Main Results. Table 2 presents the results of our training leveraging TSE in reinforcement learning across different datasets. Our observations are as follows: (1) With TSE reward alone, our method consistently improves model performance across diverse answer lengths and datasets. (2) With TSE reward alone, our method achieves performance on par with or even better than using accuracy reward, despite not relying on ground-truth answers during training. On the Countdown dataset, for example, LLaDA-8B-Instruct trained with our approach achieves a 24.7% improvement, significantly outperforming the 15.1% gain observed with d1. (3) By combining TSE with accuracy reward, our method significantly outperforms using accuracy reward alone, achieving absolute improvements of 0.9% on GSM8K, 0.2% on MATH500, 1.7% on SVAMP, and 10.2% on Countdown, compared to d1 (Zhao et al., 2025). In total, our approach yields average gains of 2.0%, 4.3%, 6.6%, and 25.3% on these benchmarks, respectively, compared to the SFT baseline. These results demonstrate the effectiveness of encouraging temporally consistent answers in post-training.
@@
Figure 6: Model attributes after reinforcement finetuning (RFT). (a) Temporal semantic entropy decreases across datasets after RFT, indicating improved semantic consistency in model outputs. (b) The ever pass rate remains higher than the final pass rate, suggesting room for continued enhancement. (c) The number of effective tokens per generation is reduced after RFT, reflecting more concise outputs. This brevity may contribute to reduced temporal oscillations, though further analysis is needed.
Model Attributes After RFT. We conduct an in-depth analysis of the LLaDA-8B-Instruct model fine-tuned with the negative TSE reward alone, with a generation length of 128 tokens, evaluating its behavior after RFT across several dimensions: temporal semantic entropy, the ever pass rate , and the number of effective tokens (defined as the average count of non-padding, non-EOS tokens per generation). Our findings reveal that (1) as shown in Fig. 6a, temporal semantic entropy consistently decreases across various datasets after RFT, reflecting enhanced temporal consistency in the model’s outputs—an anticipated result of reinforcement learning; (2) Fig. 6b demonstrates that the ever pass rate remains higher than the final pass rate after RFT, suggesting there is still potential for further improvement; and (3) as depicted in Fig. 6c, the number of effective tokens reduces following RFT, implying that the fine-tuned model generates more concise responses. We hypothesize that shorter outputs may be less prone to temporal oscillations, although this relationship requires further investigation. More analysis of the model trained after temporal consistency reinforcement is provided in the Sec. D.2.
6 Limitations
While our temporal self-consistency voting and post-training approach demonstrates effectiveness in many scenarios, it exhibits significant limitations when applied to tasks where the model’s intermediate predictions are consistently inaccurate. As discussed in Sec. D.3, for the Sudoku dataset, the average correctness across all intermediate generation steps remains exceedingly low (below 5%), making it difficult to reliably vote for the correct answer. Similarly, RFT relies on the model already achieving reasonably good performance to produce meaningful reward signals (Prabhudesai et al., 2025; Agarwal et al., 2025), though combining TSE with the accuracy reward may alleviate this issue to some extent. This underscores that our approach depends on the model’s inherent ability to generate correct or near-correct answers in the sampling trajectory.
@@ Appendix A Appendix Overview
  • Appendix B: More Implementation Details
@@
    • Sec. B.1: Detailed descriptions of the datasets used
    • Sec. B.2: Training configurations and hyperparameters
    • Sec. B.3: Sampling strategies and evaluation procedures
  • Appendix C: More Analysis
@@
    • Sec. C.1: Accuracy and entropy analysis on the MATH500 and SVAMP datasets
    • Sec. C.2: Temporal semantic entropy across varying generated lengths
    • Sec. C.3: Block-level token entropy analysis
  • Appendix D: More Experimental Results
@@
    • Sec. D.1: Ablation studies on different scoring rules for combining TSE with accuracy reward
    • Sec. D.2: Performance of temporal self-consistency voting on reinforcement fine-tuned models
    • Sec. D.3: Analysis of failure cases on the Sudoku dataset
    • Sec. D.4: Detailed examples illustrating temporal oscillation
Appendix B More Implementation Details
B.1 Datasets

### v2

@@ 4.1 Temporal Self-Consistency Voting
We propose a temporal self-consistency decoding strategy for dLLMs, which leverages intermediate predictions to improve final outputs. As discussed in Sec. 3.3, while the last timestep usually yields the best result, the correct answer may also appear earlier, so relying only on the final prediction risks discarding better outputs. To address this, we aggregate predictions across timesteps using a weighted voting mechanism. Formally, given a diffusion sampling trajectory $\left\{x_{0}^{t}\right\}_{t=1}^{T}$, our method selects the final answer $a^{*}$ according to a weighted vote over all timesteps:
@@
Here, $\mathbbm{1}(\cdot)$ indicates whether $x_{0}^{t}$ decodes to $a$, and $f(t)$ is a timestep weighting function. Since accuracy generally increases with later steps, we design $f(t)$ as a monotonically decreasing function of diffusion step. We experiment with constant, linear, and exponential weighting, as in Sec. 5.2.
Discussion. Our method is conceptually related to self-consistency decoding (Wang et al., 2022), which improves reasoning in autoregressive LLMs by sampling diverse reasoning paths and selecting the most consistent answer via majority voting. However, self-consistency requires multiple full-length forward passes with high cost. In contrast, our approach requires only a single sampling trajectory. By exploiting the temporal nature of diffusion inference indicated by Eq. 1, we obtain a series of intermediate predictions without additional model evaluations. This makes our method both efficient and effective for boosting accuracy through temporal aggregation.
@@ 4.2 Temporal Consistency Reinforcement
Negative TSE as the Reward. Following GRPO, for each question $q$ sampled from the dataset $\mathcal{D}$, we draw a group of $G$ responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. Each response $o_{i}$ receives a scalar reward $r_{i}=-\operatorname{TSE}(o_{i})$, where $\operatorname{TSE}(o_{i})$ is computed using Eq. 3 from Sec. 3.3. This reward encourages the model to produce responses whose intermediate predictions remain semantically consistent throughout the decoding process. Based on this, we define the unnormalized advantage (Liu et al., 2025b) for all tokens $k=1,\ldots,|o_{i}|$ as $A_{i}^{k}(\pi)=r_{i}(\pi)-\operatorname{mean}\left(\left\{r_{j}(\pi)\right\}_{j=1}^{G}\right).$ The training objective follows the standard GRPO formulation with our TSE-based reward:
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{\stackrel{{\scriptstyle q\sim\mathcal{D}}}{{o_{1},\ldots,o_{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}}}\left[\left(\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\min\left(\rho_{i}^{k}A_{i}^{k},\text{clip}\left(\rho_{i}^{k},1-\varepsilon,1+\varepsilon\right)A_{i}^{k}\right)\right)-\beta D_{\text{KL}}\left[\pi_{\theta}(\cdot|q)\|\pi_{\text{ref}}(\cdot|q)\right]\right],$
@@
Combining TSE with Accuracy Reward. When ground-truth answers are available during training, we combine TSE with an accuracy reward to further enhance performance. Specifically, We design a composite reward function that integrates correctness and temporal consistency.
The accuracy reward follows a binary scheme: it assigns 1 when the model’s prediction is correct ($o_{i}=o^{*}$), and 0 otherwise. To measure temporal consistency, we derive a normalized confidence score from TSE: $c(o_{i})=\frac{\mathcal{H}_{\max}-\operatorname{TSE}(o_{i})}{\mathcal{H}_{\max}}$, $\mathcal{H}_{\max}=\log T$. Here, $T$ denotes the total sampling steps. This normalization yields $c\in[0,1]$, with larger values indicating stronger consistency.
We further transform $c(o_{i})$ using the spherical scoring rule (Gneiting & Raftery, 2007), which has shown strong empirical performance across tasks. The final reward is defined as:
@@
Here, the first term enforces correctness while the second term encourages higher temporal consistency. This formulation makes the reward sensitive not only to prediction accuracy, but also to the stability of the underlying generation process. A detailed comparison of various scoring rules and their empirical performance in the Sec. F.2.
@@ 5 Experiments
Table 1: Performance of temporal majority voting. We compare three strategies: fixed, linear, and exponential weighting, on four datasets using LLaDA-8B-Instruct and LLaDA-1.5. Bold numbers mark group bests, and green values show gains over baseline. For reference, we report the oracle $\operatorname{EverPass}@1\mid t$ as an upper bound.
@@ 5.1 Implementation Details
We use LLaDA-8B-Instruct (Nie et al., 2025) and LLaDA-1.5 (Zhu et al., 2025), evaluating on four math benchmarks: GSM8K (Cobbe et al., 2021), MATH500 (Hendrycks et al., 2021), SVAMP (Patel et al., 2021), and Countdown (Pan et al., 2025). Following d1 (Zhao et al., 2025), we report performance under different output lengths. Temporal self-consistency voting applies exponential weights to sampling steps. For post-training, LLaDA-8B-Instruct undergoes supervised fine-tuning (SFT) on s1K (Muennighoff et al., 2025) for 20 epochs with 4,096 token sequences, followed by reinforcement fine-tuning (RFT). For LLaDA-1.5, we omit SFT because it often results in performance degradation, likely due to the model having already undergone sophisticated post-training. All training uses 8 H800 GPUs. Further details are in Appendix C.
@@ 5.2 Temporal Self-Consistency Voting
Voting Strategies. We apply weighted voting across denoising steps using three schemes: fixed, linear, and exponential. Each uses a weighting function $f(t)$, where $t$ is the current diffusion step. The fixed scheme assigns equal weight to all steps with $f(t)=1$. The linear weighting takes the form f(t)=1t/Tf(t)=1-t/T, and exponential uses $f(t)=\exp(\alpha(1-t/T))$ with $\alpha=5$. Both linear and exponential schemes prioritize early diffusion time steps, i.e., latter sampling steps.
Ablations on Voting Strategies. As shown in Table 1, linear and exponential weighting improve inference performance, with exponential yielding the largest gains, e.g., LLaDA-8B-Instruct improves by 1.6%, 1.2%, 1.0%, and 2.2% on GSM8K, MATH500, SVAMP, and Countdown, respectively. Fixed weighting performs slightly worse, likely because equal weights amplify inaccurate early predictions. We therefore adopt exponential weighting by default. We futher ablate the exponential hyperparameter $\alpha$. As shown in Fig. 5b, $\alpha$ values ranging from 1 to 11 consistently improve accuracy, peaking at $\alpha=5$ with an average gain of 1.5%. Thus, we set $\alpha=5$ by default.
@@ 5.3 Temporal Consistency Reinforcement
Table 2: Performance of reinforcement fine-tuning. Unlike d1 (Zhao et al., 2025), which requires ground-truth answers, our method uses negative Temporal Semantic Entropy (TSE) as the reward without labels. Combining TSE with accuracy-based rewards yields further gains across benchmarks. Green numbers denote improvements over baseline.
@@
+ RFT accuracy reward (d1) 73.0 78.9 83.1 29.8 36.2 40.2 84.7 89.3 88.0 38.7 29.7 39.1
negative TSE reward (ours) 72.0 80.8 82.6 30.2 35.0 40.0 86.3 88.3 87.3 50.0 55.9 53.1
+2.2 +1.4 +1.5 +1.2 +2.6 +4.6 +1.0 +2.0 +4.0 +28.5 +34.8 +32.4
combining both (ours) 73.2 80.5 84.0 29.6 35.4 41.4 86.3 90.3 89.0 44.5 46.9 63.3
+3.4 +1.1 +2.9 +0.6 +3.0 +6.0 +1.0 +4.0 +5.7 +23.0 +25.8 +42.6
Main Results. Table 2 reports results of incorporating temporal consistency into RFT. We have the following observations. (1) Using TSE reward alone consistently improves performance across lengths and datasets. (2) TSE reward matches or surpasses accuracy reward despite not using ground truth, e.g., on Countdown, LLaDA-8B-Instruct improves 24.7% vs. 15.1% with d1. (3) Combining TSE with accuracy reward further boosts results, with absolute gains of 0.9% (GSM8K), 0.2% (MATH500), 1.7% (SVAMP), and 10.2% (Countdown) over d1. Overall, our method achieves average improvements of 2.0%, 4.3%, 6.6%, and 25.3% over the SFT baseline, confirming the benefit of encouraging temporal consistency.
@@
Figure 6: Model attributes after reinforcement finetuning (RFT). (a) Temporal semantic entropy decreases after RFT, showing improved semantic consistency in outputs. (b) The ever pass rate remains above the final pass rate, leaving room for further gains. (c) Effective tokens per generation drop after RFT, yielding more concise outputs.
Model Attributes After RFT. We analyze LLaDA-8B-Instruct fine-tuned with the negative TSE reward, with generation length 128, evaluating its behavior across several dimensions: TSE, the ever pass rate , and the number of effective tokens (defined as the average count of non-padding, non-EOS tokens per generation). We have the following observations. (1) As shown in Fig. 6a, temporal semantic entropy consistently decreases across various datasets after RFT, reflecting enhanced temporal consistency in the model’s outputs—an anticipated result of reinforcement learning; (2) Fig. 6b demonstrates that the ever pass rate remains higher than the final pass rate after RFT, suggesting there is still potential for further improvement. (3) In Fig. 6c, effective tokens decline after RFT, implying more concise outputs. We suspect shorter generations may reduce temporal oscillations, though this requires further study. Further analysis of model attributes after RFT and a detailed discussion of limitations are provided in Sec. F.3 and Appendix E.
@@ Acknowledgment
Reproducibility Statement
We have made extensive efforts to ensure the reproducibility of our work. Detailed dataset descriptions are provided in Sec. C.1, and training configurations and hyperparameters are reported in Sec. C.2. The sampling and evaluation procedures are outlined in Sec. C.3. To further promote transparency, we provide mathematical formulations and thorough descriptions of our proposed algorithms—Temporal Self-Consistency Voting and Temporal Consistency Reinforcement—directly in the main paper. Upon acceptance, we will release our models, together with training and inference code, to facilitate replication and further research.
@@ Appendix A Appendix Overview
  • Appendix B: LLM Usage
    Clarify the assistance from large language models (LLMs).
  • Appendix C: More Implementation Details
@@
    • Sec. C.1: Detailed descriptions of the datasets used
    • Sec. C.2: Training configurations and hyperparameters
    • Sec. C.3: Sampling strategies and evaluation procedures
  • Appendix D: More Analysis
@@
    • Sec. D.1: Accuracy and entropy analysis on the MATH500 and SVAMP datasets
    • Sec. D.2: Temporal semantic entropy across varying generated lengths
    • Sec. D.3: Block-level token entropy analysis
  • Appendix E: Limitations
    Discuss limitations and analyze failure cases, including:
    • Sec. E.1: Discuss potential limitations of our method
    • Sec. E.2: Analysis of failure cases on the Sudoku dataset
  • Appendix F: More Experimental Results
@@
    • Sec. F.1: Training a unified model across multiple tasks
    • Sec. F.2: Ablation studies on different scoring rules for combining TSE with accuracy reward
    • Sec. F.3: Performance of temporal self-consistency voting on reinforcement fine-tuned models
    • Sec. F.4: Detailed examples illustrating temporal oscillation
  • Appendix G: Change Log
    Summarizes the main modifications introduced in arXiv v2 compared to v1.
Appendix B LLM Usage
In this section, we clarify the role of large language models (LLMs) in preparing this work. The model was used exclusively for language polishing, such as refining grammar, style, and readability, without contributing to the research design, analysis, or conclusions.
Appendix C More Implementation Details
C.1 Datasets